10x+6y=38 and 20x+4y=52 please solve and provide work

Answer by macston(5194)

10x+6y=38 Multiply both sides by 2
20x+12y=76
20x+4y=52 Subtract this from above.
8y=24 Divide each side by 8.
y=3 Plug value into one of the original equations and solve for x.
10x+6y=38
10x+6(3)=38
10x+18=38 Subtract 18 from each side.
10x=20 Divide each side by 10.
x=2 ANSWER:(2,3)
CHECK:
Use the other original equation to check.
20x+4y=52
20(2)+4(3)=52
40+12=52
52=52